OpenStudy (anonymous):

Please help me... award given if solved solve the system using the substitution method x+y=4 2x+3y=-2

OpenStudy (anonymous):

-2x+-2y=-8 2x+3y=-2 cancels out the 2x -5y=-10 y=2 plug y=2 into first equation: x+2=4 x=2 (2,2) is the final answer